OpenStudy (anonymous):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Show your work. Sn: 2 is a factor of n2 + 7n

OpenStudy (chealyn98):

A statement Sn about the positive integers is given. Write statements S1, S2, and S3, and show that each of these statements is true. Show your work. Sn: 1^2 + 4^2 + 7^2 + . . . + (3n - 2)^2 = (n(〖6n〗^2-3n-1))/2 S1:1^2+4^2+7^2+…+(3*1-2)^2=1(6*1^2-3n-1)/2 S1: 1^2...